Mr. Hoosier has gathered all these data about his finances.

Salary

$159,000

Taxable interest

$5600

Municipal Bond interest

$20,000

Total Itemized deductions

$9500

The personal exemption is $4700. The standardized deduction for a single filer is 5000.

Single Filer

If Taxable Income is

Then tax is:

0-8000

10% of the amount over 0

8000- 35,000

800 + 15% of the amount over 8000

35000- 55,000

4850 + 20% of the amount over 35000

55000- 150,000

8850 + 30% of the amount over 55000

150000 – 200000

37350 + 35% of the amount over 150000

200000 and over

54850 + 40% of the amount over 200000

Married filing jointly:

If Taxable Income is

Then tax is:

0-8000

10% of the amount over 0

8000- 35,000

800 + 18% of the amount over 8000

35000- 55,000

4850 + 22% of the amount over 35000

55000- 150,000

8850 + 35% of the amount over 55000

150000 – 200000

37350 + 37% of the amount over 150000

200000 and over

54850 + 45% of the amount over 200000

a. What is his tax?

b. His average effective tax rate

c. His average tax rate

d. His marginal tax rate

e. Assume he discovers that he is eligible for a 500 tax deduction. How much does his tax liability fall with

that addition?

f. Now he discovers that he is eligible for a $500 tax credit. How much does his tax liability fall due to this

credit?
Answer #1

How is the Tax Calculation Done?

In given case, if you are a filing your taxes as 'single', have no children, are below the age of 65, and you are not blind, with the consideration of the standard tax deduction will be:

Calculations based on preliminary 2019 tax tables:

Tax computation

Gross income

USD

Salary

159,000

Interest

5600

Municipal Bond interest

Exempt

Total gross income

164600

Adjustments

0

Adjusted gross income (AGI)

164600

Less - standard deduction

50 00

Taxable income

159600

Tax thereon

On taxable income of 159600

40710

(37350+35% of (159600-150000))

Total tax before credits

40710

Less - credits

Net tax

40710

  1. His net tax $ 40710
  2. His average Effective tax rate

= 40710/159600

= 25.51%

  1. His average tax rate

= 25.51%

  1. Marginal tax rate

= 35%

  1. IF he is eligible for a 500 tax deduction then his tax liability –

=Net tax – Tax deduction

= 40710-500

= $40210

  1. IF he is eligible for a 500 tax credit then his tax liability –

=Total tax before credit – Credit

= 40710-500

= $40210